MsBrains (ms-brains):

EASY! PLEASE HELP! will medal and fan! 2(3(10) - -3) solve step by step I know the answer is 66, but how?

OpenStudy (anonymous):

What's 3 times 10?

MsBrains (ms-brains):

30, @alina1234

OpenStudy (dayakar):

2[3(10)-(-3)] is this ur quastion

OpenStudy (dayakar):

first multiply 3*10 =30 _(-3) = +3 now it becomes like this 2[30+3]

OpenStudy (dayakar):

=2*33 =66

OpenStudy (dayakar):

is it clear

MsBrains (ms-brains):

Wait, why did -3 become positive?

MsBrains (ms-brains):

@dayakar

OpenStudy (johan14th):

-3 became positive because the minus sign between 10 and 3 and the negative sign in front of 3 cancel each other out and make a positive.

OpenStudy (johan14th):